Current alignment light status
All Lights operating and in place.  Waiting on software for 
remote and automatic operation

If Short delay appears to want to be at a strange location check that
the offset is reasonable.  NAN indicates that the fringe tracker code
did not know what to put in and loaded garbage.  Clear with
delay paddle>sdx>clear fringe tracker offsets

Short delay laser takes 4 minutes to warm up and you should not
attempt to do anything with the SD system until that condition clears.
Also, the two enable lines are tied together so you need to
turn SD1 on and SD2 on before you issue a command like SD2 home.

Long delay lines are initialized to zero whenever 
"server" is started.  Homing seems fine these days.
In the distant past is sometimes overshot the end of travel and hit the
mechanical limit which required manual intervention
(disconnecting the 5V line which is currently jumpered on
the CB Com in the laser hut)

Long delay seems to dislike having commands stacked - 
system crashed a few times after LD2 motion was initiated before
LD1 motion was completed.  Correction was reboot of the whole system.
It is possible that unloading/reloading ./srv2  task would fix it but that
was not tried.  It may the the lower level one which becomes corrupt.
PS  Someone else tried to have both axes moving we had a problem again.
This time we cycled power on the AT6200 and then reloaded ./srv2 as
root and the system worked.  There is an operating system download
(done at bootup) which should be missing with this procedure but things 
seemed to work.  If we crash again I would like to try just
the reload of the server code.

After cycling power  on the VME the programs are not loaded and you need
to reload with
vxworks-start-LL and daq-start
Telescope locations will be forgotten so if you note where they were
when the system hung and move them to the negative then you will be
close to home when you issue the home command.  or you can move
to the acutal location with the telescopes powered down.
Do not  home the telescopes!

Star tracker chiller takes significant current.
The triplite plug strip circuit breaker tripped once when
we had the LList, 4 monitors, 3 oscilloscopes, 3 more monitors 
all on.  Need to get another circuit in.

Flip in LED's are controlled by small switches.  The switches have
been known to fail and stick either on or off.  Sometimes mechanical
intervention is required or replacement of the switch.

Printing queue on iota16  iota 17 can be blocked if an A4 page is 
sent to the printer.
to clear
lpq to see print job ID's
lprm to remove the offending job

Marc Lacasse